OpenStudy (anonymous):

how do I simplify the square root of 98x^3

OpenStudy (anonymous):

3 square root 10 im pretty sure....? do a factor tree to get an answer. you can do it.

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ 98\implies 2\cdot 7\cdot 7\implies 2\cdot 7^2\qquad \qquad x^3\implies x^2\cdot x}} \\ \quad \\ \quad \\ \sqrt{98x^3}\implies \sqrt{2\cdot 7^2\cdot x^2\cdot x}\) what do you think?